While driving on the highway Fiona averages 60 miles per hour. In the city she averages 30 miles per hour. On a trip to her grandmother's house Fiona drove 312 hours on the highway and 112 hours in the city.
What was Fiona's average speed for the entire trip?


Answer by TimothyLamb(4379)

s = d/t
d = st
---
x = distance on highway
y = distance on city streets
---
distance on highway
x = 60*312
x = 18720
---
distance on city streets
y = 30*112
y = 3360
---
total distance:
x + y = 18720 + 3360
x + y = 22080 miles
---
total time:
312 + 112 = 424 hours
---
average speed for entire trip:
s = 22080/424
s = 52.1 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php